OpenStudy (anonymous):

4y+5x-16=0 SOLVE A LINEAR FUNCTION !!

OpenStudy (anonymous):

so you mean like 4y=16-5x thus y=4-(5/4)x ?

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

I have to solve for y ??

OpenStudy (anonymous):

the equation would equal to y=5/4x-4

OpenStudy (anonymous):

You forgot that you need to substract 5x-16 from both sides not add them. Thus y=4-(5/4)x. @Emily55555 the thing is that 4y+5x-16=0 Is the equation of linear function so i dunno what it is that we are suppoused to do here.

OpenStudy (anonymous):

Are you trying to find the solutions?